Circular No. 6264

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

HT CASSIOPEIAE J. A. Mattei, AAVSO, reports that this SU UMa-type dwarf nova

is in outburst for the first time since 1987 Feb., its previous quiesence durations having been much shorter. Because the Hubble Space Telescope is scheduled to observe HT Cas on Nov. 23, she encourages complementary photometric and spectroscopic observations. Selected visual magnitudes: Nov. 12.08 UT, [16.0 (D. York, Abiquiui, NM); 16.09, 15.8 (York); 17.733, 13.2 (P. Schmeer, Bischmisheim, Germany); 17.912, 13.4 (W. Worraker, Didcot, U.K.); 17.99, 13.4 (R. Stewart, Fairlawn, NJ); 18.02, 13.4 (J. Bortle, Stormville, NY); 18.30, 13.1 (J. Griesé, Stamford, CT).

2EG J0432+2910 W. Voges and G. Kanbach, Max-Planck-Institut für Extraterrestrische

Physik, Garching, report: “Near the radio position of 2EG J0432+2910 (IAUC 6258), we have found a hard x-ray source, RX J0433.6+2906, in the ROSAT all-sky survey data centered on α = 4^{h}33^{m}39^{s}, δ = +29^{o}06′.2 (equinox 2000.0; estimated error 25′′ in 1σ radius), which is within 22′′ of the radio position; see also Neumann et al. (1994, A.Ap. Suppl. 106, 303), who refer to the object as EF 0430+2859. The PSPC count rate in the ROSAT band (0.1–2.4 keV) was 0.05 count/s during 1990 Aug. 21–23. The energy flux is ∼ 2 × 10^{−12} erg cm^{−2} s^{−1}. This is the only ROSAT x-ray source within a radius of 1^{o}. Condon et al. (1983, A.J. 88, 20) quote flux densities of 0.22, 0.37, and 0.36 Jy at 1.4, 2.695, and 8.085 GHz, respectively, from interferometer observations made in 1978 of a source 0430+290, which they optically identified with an extended (∼ 5′′) galaxy (estimated mag 19.0).”

COMET 15P/FINLAY Total CCD magnitudes from J. V. Scotti, Kitt Peak (0.9-m Spacewatch

telescope): Oct. 26.47 UT, 20.2; Nov. 18.52, 20.5.

COMET 45P/HONDA-MRKOS-PAJDUSÁKOVÁ Total CCD magnitudes: Oct. 21.21 UT, 21 (D. J. Tholen, Mauna

Kea, 2.24-m reflector); 22.21, 21 (Tholen); Nov. 16.38, 15.9 (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector); 18.38, 15.7 (Nakamura).

1995 November 18 Daniel W. E. Green